# Unsupervised cell segmentation by fast Gaussian Processes

### v1

@@ 1 Introduction
Cellular segmentation tools can be broadly classified as unsupervised and supervised methods, based on whether labelled data is needed for training the models. One of the most popular tools of unsupervised cellular segmentation is the ImageJ abramoff2004image ; schneider2012nih ; schindelin2012fiji , which enables segmenting both cell nuclei and whole cells, and linking cells across frames of a microscopy video via plug-ins such as TrackMate tinevez2017trackmate . Prior to using ImageJ for cell detection and segmentation, a series of user-driven image pre-processing steps must be completed in the program. For instance, the image is typically converted into a binary image first, based on the threshold chosen either manually or automatically via several global threshold setting methods nichele2020quantitative ; liu2009otsu . Additional image processing steps, such as despeckling for noise reduction ferreira2011imagej , typically need to be run before segmentation. To detect cell objects within an image, the algorithm scans the corresponding binary image until the edge of an object is reached. Then, the algorithm traces around the edge of the object until the starting point is reached. This process is repeated for pixels of each object, or the foreground pixels, in the binary image. The centroid of each cell object can be identified by fitting an ellipse for each object using the second-order central moments schneider2012nih , which assumes the cell boundary is elliptical. The parameters for the best-fit ellipse, like the major and minor axes and centroids, for each object boundary are returned, with the option to return matrices for the object masks and non-elliptical outlines as well ferreira2011imagej . Additional functionality can be added to the user interface by writing macro code to streamline image analysis or adding third-party plug-ins schneider2012nih . However, achieving reasonable segmentation results is often labor-intensive and requires human interaction at each step. Furthermore, The assumption of the shape of the cells in ImageJ is restrictive. Though the ellipse can approximate nuclei of certain cells, it cannot approximate the shape of cytoplasm or the whole cell of frequently used cells in experiments, such as fibroblasts or epithelial cells.
@@
Supervised cellular segmentation methods, are typically built upon a convolutional neural network structure, such as the U-net ronneberger2015u . A popular tool is $\sf{Cellpose}$, which builds U-nets trained on a dataset of over 70,000 segmented objects, including annotated images that contain numerous types of cells stringer2021cellpose . The model first compresses the image and then reconstructs the image to its original dimensions through a U-net to capture both local and global features. It further generates a topological map from diffusion simulation to product gradient maps for the model to learning the horizontal and vertical gradients of the image. The final layer of Cellpose produces three feature maps, including the horizontal and vertical directional gradients, and the sigmoid transformation of the probability map of the output image from the U-net, and the parameters of the Cellpose are then estimated by fitting a $L_{2}$ loss function of the gradient vector and binary label of each pixel of the image. As the Cellpose segmentation requires massive training data, it may not work well for detecting novel cell types, where high-labeled image data that may not be readily available. Furthermore, training Cellpose and using it for segmenting large images with several million pixels can be computationally expensive.
@@
In response to the challenges the current methods face, we present a novel unsupervised cell image segmentation method based on separable Gaussian processes, automated thresholding, and watershed operations for image segmentation. The proposed method has several advantages. First, the separable covariance structure on image data enables us to decompose the large covariance into a Kronecker product of two small covariances, which substantially reduces the computational cost for computing the matrix inversion and the logarithm of the determinant of the covariance matrix. Second, unlike conventional unsupervised image segmentation methods such as ImageJ, which require manual tuning of multiple hyperparameters, including Gaussian blur radius, thresholding method, and minimum object size, our proposed segmentation pipeline is free of manual parameter selection. We developed fast algorithms for maximum likelihood estimation of the parameters in GPs and the predictive distribution. The predictive mean surface naturally facilitates foreground-background separation by applying a data-driven threshold with the optimal threshold automatically estimated based on absolute second difference in foreground pixel counts between the thresholds. Thrid, the cell masks are segmented using the watershed algorithm, which simulates water flowing through a topographical map of the image to separate cell objects. Fourth, our approach is entirely unsupervised, negating any need for annotated training data, in contrast to the supervised methods, such as Cellpose. Thus, the proposed method is particularly suitable for segmenting new cell types or other biological structures. We studied our methodology by simulated data for noise filtering. We also compared our approaches with the alternatives using real microscopy images for segmenting both the cell nucleus and whole cell in optically dense images, a challenging scenario even when labelled data are available. These studies demonstrate that our approach substantially improves the alternatives.
@@ 2.1 The likelihood function and predictive distribution
where $\mathbf{F}$ follows a matrix normal distribution $\mathbf{F}\sim\mbox{Matrix-Normal}\left(\mu\mathbf{1}_{n_{1}\times n_{2}},\,\sigma^{2}\mathbf{R}_{1},\mathbf{R}_{2}\right)$ with $\mu$ being a mean parameter, $\mathbf{1}_{n_{1}\times n_{2}}$ being a $n_{1}\times n_{2}$ matrix of ones, $\sigma^{2}$ being the variance parameter, $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$ being correlation matrices for two inputs, respectively, and the noise matrix follows $\mathbf{E}\sim\mathcal{MN}(0,\sigma^{2}_{0}\mathbf{I}_{N\times N})$, with $N=n_{1}n_{2}$. Additional trend structure can be modeled in the mean function, and the covariance can be generalized to be semi-separable in the model gu2022gaussian .
@@
$(\mathbf{y}_{v}\mid\sigma^{2},\eta,\mathbf{R}_{1},\mathbf{R}_{2})\sim\mathcal{MN}(\mathbf{\mu}\mathbf{1}_{N},\sigma^{2}\left(\mathbf{R}_{2}\otimes\mathbf{R}_{1}+\eta\mathbf{I}_{N\times N})\right),$ (2)
where $\otimes$ is a Kronecker product and $\eta=\sigma^{2}_{0}/\sigma^{2}$ is a nugget parameter. Here the correlation matrices are parameterized kernel functions $R_{1}(i,i^{\prime})=K_{1}(x_{i,1},x_{i^{\prime},1})$ and $R_{2}(j,j^{\prime})=K_{2}(x_{j,2},x_{j^{\prime},2})$. Frequently used kernel functions include the power exponential kernel and Matérn kernel rasmussen2006gaussian . Denote $K_{l}(x_{i,l},x_{i^{\prime},l})=K_{l}(d_{l})$ with $d_{l}=|x_{i,l}-x_{i^{\prime},l}|$. For instance, the Matérn kernel with roughness parameter $5/2$ follows handcock1993bayesian :
@@
where $\gamma_{l}$ is a range parameter to be estimated for $l=1,2$. A GP with the kernel function in Equation (3) is twice mean-squared differentiable rasmussen2006gaussian , and it is used as a default choice of some software packages of GP surrogate models roustant2012dicekriging ; gu2018robustgasp . We use the kernel function in (3) for demonstration purposes. where $\gamma_{l}$ is a range parameter to be estimated for $l=1,2$. A GP with the kernel function in Equation (3) is twice mean-squared differentiable rasmussen2006gaussian , and it is used as a default choice of some software packages of GP surrogate models roustant2012dicekriging ; gu2018robustgasp . We use the kernel function in (3) for demonstration purposes.
where $\gamma_{l}$ is a range parameter to be estimated for $l=1,2$. A GP with the kernel function in Equation (3) is twice mean-squared differentiable rasmussen2006gaussian , and it is used as a default choice of some software packages of GP surrogate models roustant2012dicekriging ; gu2018robustgasp . We use the kernel function in Equation (3) for demonstration purposes. After specifying the model, we compute the likelihood function and predictive distribution for parameter estimation and prediction, respectively. The parameters of the model in (2) include the mean, variance, range and nugget parameters $\{\mu,\sigma^{2},\bm{\gamma},\eta\}$ with $\bm{\gamma}=(\gamma_{1},\gamma_{2})$ being the range parameters in the kernel function in Equation (3). We employed the maximum likelihood estimator (MLE) to estimate the parameters. Given the range and nugget parameters, the MLE of the mean and variance parameters has a closed-form expression: $\hat{\mu}=(\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{1}_{N})^{-1}\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{y}_{v}$, with $\mathbf{\tilde{R}}=\mathbf{R}+\eta\mathbf{I}_{N}$, and $\hat{\sigma}^{2}=\frac{S^{2}}{N}$, with $S^{2}=(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})^{T}\bm{\tilde{}}{\mathbf{R}}^{-1}(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})$. The logarithm of the profile likelihood after plugging the MLE of the mean and variance parameters follows:
@@
$(\bm{\hat{\gamma}},\hat{\eta})=\operatorname*{argmax}_{\bm{\gamma},\eta}\log(L(\bm{\gamma},\eta)).$ (5)
Vectorize signal $\mathbf{f}_{v}=\mbox{vec}(\mathbf{F})$ from model (1). After plugging the MLE of the parameters are, the posterior distribution of the image vector follows follows
@@ 2.2 Fast computation for images
The log profile likelihood in Equation (4) and the predictive distributions in Equation (6) can be written as a function in terms of eigenpairs in Lemma 1 below, where the proof is provided in the Appendix.
@@
    $\log(L(\bm{\gamma},\eta))=C-\frac{1}{2}\sum^{n_{1}}_{i=1}\sum^{n_{2}}_{j=1}\log\left(\lambda_{i,1}\lambda_{j,2}+\eta\right)-\frac{N}{2}\log\left(\sum^{n_{1}}_{i=1}\sum^{n_{2}}_{j=1}\frac{\tilde{Y}_{i,j}^{2}}{\lambda_{i,1}\lambda_{j,2}+\eta}\right),$ (9)
@@
    $\hat{\mu}=\left(\sum^{n_{2}}_{j=1}\sum^{n_{1}}_{i=1}\frac{\tilde{u}^{2}_{i,1}\tilde{u}^{2}_{j,2}}{\lambda_{i,1}\lambda_{j,2}+\eta}\right)^{-1}\left(\sum^{n_{2}}_{j=1}\sum^{n_{1}}_{i=1}\frac{\tilde{u}_{i,1}\tilde{Y}_{i,j,0}\tilde{u}_{j,2}}{\lambda_{i,1}\lambda_{j,2}+\eta}\right),$ (10)
@@ 3 Image segmentation from Gaussian processes
Figure 1: Workflow for segmenting and labeling cell images: (A) Divide the image into different sub-images to enable locally estimated mean and variance parameters for capturing local properties such as change of brightness. (B) Compute the predictive mean of fast GPs in Section 2.2 to each sub-image, which greatly reduces image noise. (C) Threshold each smoothed sub-image based on the criterion discussed in Section 3.2 to produce binary images, separating cells from the background. The optimal threshold is estimated for each sub-image. (D) Recombine the binary sub-images into a single binary image, and apply the watershed algorithm discussed in Section 3.3 to the image for segmentation and labeling, resulting in color-coded cell masks.
@@
Though Gaussian processes have been widely used in as surrogate models to emulate expensive computer simulations, nonparametric regression, and modeling spatially correlated data rasmussen2006gaussian ; chang2016calibrating , the flexibility and scalability of GPs in unsupervised image segmentation have not been studied yet. The overall flow of our cellular image segmentation process is summarized in Figure 1. First, some microscopy images of cells are particularly large with $10^{6}-10^{7}$ pixels. Similar to other image segmentation methods schneider2012nih ; stringer2021cellpose , we segment each large image into multiple sub-images. We then apply the fast algorithm of GPs in Lemma 1 to get the predictive mean of each sub-image, illustrated in Parts (A) and (B) in Figure 1, respectively. We assume the range parameters and nugget of the GPs are the same for each subimage estimated by MLE in Lemma 1, yet the mean and variance parameters are distinct for each separable image, and the estimation of these parameters has a closed-form expression. Such a step enables different estimated parameters to capture the change of optical properties, such as brightness difference and out-of-focus blur from distinct parts of a large microscopy image, which are crucial in practice. Second, we developed a robust measure that will be introduced in Section 3.2 to threshold the predictive mean of each sub-image into background and objects, showed in Panel (C). Finally, the binary submatrices are restitched in Panel (D), and the watershed algorithm is applied to the binary matrix to retrieve the labeled cell masks, which will be introduced in Section 3.3.
3.1 Predictive mean from Gaussian processes on lattice
@@
Let $\mathbf{Y}$ represent the original cell image, which has size $N=n_{1}\times n_{2}$. We denote $y(\mathbf{x}_{i,j})$ as the value of the image at pixel $\mathbf{x}_{i,j}=(x_{i,1},x_{j,2})$. Microscopy images of cells can contain more than $10^{6}$ pixels. For a large image $\mathbf{Y}$, it is normally partitioned into a set of cropped images, denoted as $\{\mathbf{Y}_{k}\}_{k=1}^{\tilde{K}}$, where $\tilde{K}$ is the total number of cropped images as illustrated in Figure 1.
@@
For each cropped image, we first use GPs with fast computation introduced in Section 2.1 for smoothing the images. The range and nugget parameters of GPs are estimated by maximizing the likelihood in Equation (9), and they are held the same for all sub-images as they have similar smoothness properties. The mean and variance parameters are estimated differently for each cropped image to capture the local change of the optical properties. As the sub-images will be turned into a binary image, the discontinuity of the boundary between different sub-images does not impact the image segmentation task.
3.2 Binary cell masks

### v2

@@ 1 Introduction
Cellular segmentation tools can be broadly classified as unsupervised and supervised methods, based on whether labeled data is needed for training the models. One of the most popular tools of unsupervised cellular segmentation is the ImageJ abramoff2004image ; schneider2012nih ; schindelin2012fiji , which enables segmenting both cell nuclei and whole cells, and linking cells across frames of a microscopy video via plug-ins such as TrackMate tinevez2017trackmate . Prior to using ImageJ for cell detection and segmentation, a series of user-driven image pre-processing steps must be completed in the program. For instance, the image is typically converted into a binary image first, based on the threshold chosen either manually or automatically via several global threshold setting methods nichele2020quantitative ; liu2009otsu . Additional image processing steps, such as despeckling for noise reduction ferreira2011imagej , typically need to be run before segmentation. To detect cell objects within an image, the algorithm scans the corresponding binary image until the edge of an object is reached. Then, the algorithm traces around the edge of the object until the starting point is reached. This process is repeated for pixels of each object, or the foreground pixels, in the binary image. The centroid of each cell object can be identified by fitting an ellipse for each object using the second-order central moments schneider2012nih , which assumes the cell boundary is elliptical. The parameters for the best-fit ellipse, including the major and minor axes and centroids, and matrices for the object masks for each object boundary, are recorded in ImageJ ferreira2011imagej . Additional functionality, such as extension of elliptical outlines, can be added to the user interface by writing macro code to streamline image analysis or adding third-party plug-ins schneider2012nih . However, achieving reasonable segmentation results is often labor-intensive and requires human interaction at each step. Furthermore, the assumption of the shape of the cells in ImageJ is restrictive. Though the ellipse can approximate nuclei of certain cells, it cannot approximate the shape of cytoplasm or the whole cell of frequently used cells in experiments, such as fibroblasts or epithelial cells.
@@
Supervised cellular segmentation methods, are typically built upon a convolutional neural network structure, such as the U-net ronneberger2015u . A popular tool is $\sf{Cellpose}$, which builds U-nets for segmenting images that contain one type of cells that are trained on a dataset of over 70,000 segmented objects, including annotated images that contain numerous types of cells stringer2021cellpose . The model first compresses the image and then reconstructs the image to its original dimensions through a U-net to capture both local and global features. It further generates a topological map from diffusion simulation to product gradient maps for the model to learn the horizontal and vertical gradients of the image. The final layer of Cellpose produces three feature maps, including the horizontal and vertical directional gradients, and the sigmoid transformation of the probability map of the output image from the U-net, and the parameters of the Cellpose are then estimated by fitting a loss function of the gradient vector and binary label of each pixel of the image. As the Cellpose segmentation requires massive training data, it may not work well for detecting novel cell types, where high-labeled image data that may not be readily available. Furthermore, training Cellpose and using it for segmenting large images with several million pixels can be computationally expensive.
@@
In response to the challenges the current methods face, we present a novel unsupervised cell image segmentation method for segmenting images containing one type of cells, based on separable Gaussian processes, automated thresholding, and watershed operations for image segmentation. The proposed method has several advantages. First, the separable covariance structure on image data enables us to decompose the large covariance into a Kronecker product of two small covariances, which substantially reduces the computational cost for computing the matrix inversion and the logarithm of the determinant of the covariance matrix. Second, unlike conventional unsupervised image segmentation methods such as ImageJ, which require manual tuning of multiple hyperparameters, including Gaussian blur radius, thresholding method, and minimum object size, our proposed segmentation pipeline is free of manual parameter selection. We developed fast algorithms for maximum likelihood estimation of the parameters in GPs and the predictive distribution. The predictive mean surface naturally facilitates foreground-background separation by applying a data-driven threshold with the optimal threshold automatically estimated based on the absolute second difference in foreground pixel counts between the thresholds. Third, the cell masks are segmented using the watershed algorithm, which simulates water flowing through a topographical map of the image to separate cell objects. Fourth, our approach is entirely unsupervised, negating any need for annotated training data, in contrast to the supervised methods, such as Cellpose. Thus, the proposed method is particularly suitable for segmenting new cell types or other biological structures. We studied our methodology by simulated data for noise filtering. We also compared our approaches with the alternatives using real microscopy images for segmenting both the cell nucleus and whole cell in optically dense images, a challenging scenario even when labeled data are available. These studies demonstrate that our approach substantially improves the alternatives.
@@ 2.1 The likelihood function and predictive distribution
where $\mathbf{F}$ follows a matrix normal distribution $\mathbf{F}\sim\mbox{Matrix-Normal}\left(\mu\mathbf{1}_{n_{1}\times n_{2}},\,\sigma^{2}\mathbf{R}_{1},\mathbf{R}_{2}\right)$ with $\mu$ being a mean parameter, $\mathbf{1}_{n_{1}\times n_{2}}$ being a $n_{1}\times n_{2}$ matrix of ones, $\sigma^{2}$ being the variance parameter, $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$ being correlation matrices for two inputs, respectively, and the vectorized noise matrix follows $\mbox{Vec}(\mathbf{E})\sim\mathcal{MN}(0,\sigma^{2}_{0}\mathbf{I}_{N\times N})$, with $N=n_{1}n_{2}$. Additional trend structure can be modeled in the mean function, and the covariance can be generalized to be semi-separable in the model gu2022gaussian .
@@
${(\mathbf{y}_{v}\mid\sigma^{2},\eta,\mathbf{R}_{1},\mathbf{R}_{2})\sim\mathcal{MN}(\mathbf{\mu}\mathbf{1}_{N},\sigma^{2}\mathbf{\tilde{R}}),}$ (2)
where $\mathbf{\tilde{R}}=\mathbf{R}+\eta\mathbf{I}_{N}$, $\mathbf{R}=\mathbf{R}_{2}\otimes\mathbf{R}_{1}$, $\otimes$ is a Kronecker product and $\eta=\sigma^{2}_{0}/\sigma^{2}$ is a nugget parameter. Here the correlation matrices are parameterized kernel functions $R_{1}(i,i^{\prime})=K_{1}(x_{i,1},x_{i^{\prime},1})$ and $R_{2}(j,j^{\prime})=K_{2}(x_{j,2},x_{j^{\prime},2})$. Frequently used kernel functions include the power exponential kernel and Matérn kernel rasmussen2006gaussian . Denote $K_{l}(x_{i,l},x_{i^{\prime},l})=K_{l}(d_{l})$ with $d_{l}=|x_{i,l}-x_{i^{\prime},l}|$. For instance, the Matérn kernel with roughness parameter $5/2$ follows handcock1993bayesian :
@@
where $\gamma_{l}$ is a range parameter to be estimated for $l=1,2$. A GP with the kernel function in Equation (3) is twice mean-squared differentiable rasmussen2006gaussian , and it is used as a default choice of some software packages of GP surrogate models roustant2012dicekriging ; gu2018robustgasp . We use the kernel function in (3) for demonstration purposes.
After specifying the model, we compute the likelihood function and predictive distribution for parameter estimation and prediction, respectively. The parameters of the model in (2) include the mean, variance, range and nugget parameters $\{\mu,\sigma^{2},\bm{\gamma},\eta\}$ with $\bm{\gamma}=(\gamma_{1},\gamma_{2})$ being the range parameters in the kernel function in Equation (3). As the number of pixels is large, we employed the maximum likelihood estimator (MLE) to estimate the parameters, as the result is similar to the robust estimation by marginal posterior mode Gu2018robustness . Given the range and nugget parameters, the MLE of the mean and variance parameters has a closed-form expression: $\hat{\mu}=(\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{1}_{N})^{-1}\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{y}_{v}$, where $\hat{\sigma}^{2}=\frac{S^{2}}{N}$ with $S^{2}=(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})^{T}\bm{\tilde{}}{\mathbf{R}}^{-1}(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})$. The logarithm of the profile likelihood after plugging the MLE of the mean and variance parameters follows:
@@
$(\bm{\hat{\gamma}},\hat{\eta})=\operatorname*{argmax}_{(\bm{\gamma},\eta)}\log(L(\bm{\gamma},\eta)).$ (5)
Vectorize signal $\mathbf{f}_{v}=\mbox{vec}(\mathbf{F})$ from model (1). After plugging the MLE of the parameters, the posterior distribution of the image vector follows follows
@@ 2.2 Fast computation for images
The log profile likelihood in Equation (4) and the predictive distributions in Equation (6) can be written as a function in terms of eigenpairs in Lemma 1 below. Lemma 1 eliminates the need to invert the $N\times N$ covariance matrix and reduce the computational complexity to $\mathcal{O}(n_{1}^{3}+n_{2}^{3})$, which is significantly smaller than $\mathcal{O}(n_{1}^{3}n_{2}^{3})$. The proof of Lemma 1 is provided in the Appendix.
@@
    $\log(L(\bm{\gamma},\eta))=$ $C-\frac{1}{2}\sum^{n_{1}}_{i=1}\sum^{n_{2}}_{j=1}\log\left(\lambda_{i,1}\lambda_{j,2}+\eta\right)$-
    $\frac{N}{2}\log\left(\sum^{n_{1}}_{i=1}\sum^{n_{2}}_{j=1}\frac{\tilde{Y}_{i,j}^{2}}{\lambda_{i,1}\lambda_{j,2}+\eta}\right),$ (9)
@@
    $\hat{\mu}$= $\left(\sum^{n_{2}}_{j=1}\sum^{n_{1}}_{i=1}\frac{\tilde{u}^{2}_{i,1}\tilde{u}^{2}_{j,2}}{\lambda_{i,1}\lambda_{j,2}+\eta}\right)^{-1}\times$
    $\left(\sum^{n_{2}}_{j=1}\sum^{n_{1}}_{i=1}\frac{\tilde{u}_{i,1}\tilde{Y}_{i,j,0}\tilde{u}_{j,2}}{\lambda_{i,1}\lambda_{j,2}+\eta}\right),$ (10)
@@ 3 Image segmentation from Gaussian processes
Figure 1: Workflow for segmenting and labeling cell images: (A) Divide the image into different sub-images to enable locally estimated mean and variance parameters for capturing local properties such as change of brightness. (B) Compute the predictive mean of fast GPs in Section 2.2 to each sub-image, which greatly reduces image noise. (C) Threshold each smoothed sub-image based on the criterion discussed in Section 3.2 to produce binary images, separating cells from the background. The optimal threshold is estimated for each sub-image. (D) Recombine the binary sub-images into a single binary image, and apply the watershed algorithm discussed in Section 3.3 to the image for segmentation and labeling, with each cell marked by a unique color.
@@
Though Gaussian processes have been widely used in as surrogate models to emulate expensive computer simulations, nonparametric regression, and modeling spatially correlated data rasmussen2006gaussian ; chang2016calibrating , the flexibility and scalability of GPs in unsupervised image segmentation have not been studied yet. The overall flow of our cellular image segmentation process is summarized in Figure 1. First, some microscopy images of cells are particularly large with $10^{6}-10^{7}$ pixels. Similar to other image segmentation methods schneider2012nih ; stringer2021cellpose , we segment each large image into multiple sub-images. We then apply the fast algorithm of GPs in Lemma 1 to get the predictive mean of each sub-image, illustrated in Parts (A) and (B) in Figure 1, respectively. We assume the range parameters and nugget of the GPs are the same for each subimage estimated by MLE in Lemma 1, yet the mean and variance parameters are distinct for each separable image, and the estimation of these parameters has a closed-form expression. Such a step enables different estimated parameters to capture the change of optical properties, such as brightness difference and out-of-focus blur from distinct parts of a large microscopy image, which are crucial in practice. Second, we developed a robust measure that will be introduced in Section 3.2 to threshold the predictive mean of each sub-image into background and objects, showed in Panel (C). Finally, the binary sub-images are restitched in Panel (D), and the watershed algorithm is applied to the binary matrix to retrieve the labeled cell masks, which will be introduced in Section 3.3.
3.1 Image denoising from Gaussian processes on lattice
@@
Let $\mathbf{Y}$ represent the original cell image, which has size $N=n_{1}\times n_{2}$. We denote $y(\mathbf{x}_{i,j})$ as the value of the image at pixel $\mathbf{x}_{i,j}=(x_{i,1},x_{j,2})$. All possible pixel values are between 0 and 1, as each image in this analysis is in grayscale. Microscopy images of cells can contain more than $10^{6}$ pixels, and some areas of an image can have substantially higher pixel values than other areas. Thus, for a large image $\mathbf{Y}$, it is normally partitioned into a set of cropped images, denoted as $\{\mathbf{Y}_{k}\}_{k=1}^{\tilde{K}}$, where $\tilde{K}$ is the total number of cropped images as illustrated in Figure 1. This step can substantially reduce the computational cost, and enable segmentation outcomes more adaptive to the features of the local areas in a large image.
@@
As experimental images often contain substantial noise, for each cropped image, we first use GPs with fast computation introduced in Section 2.1 for denoising the images. The range and nugget parameters of GPs are estimated by maximizing the likelihood in Equation (9), and they are held the same for all sub-images as they have similar smoothness properties. The mean and variance parameters are estimated differently for each cropped image to capture the local change of the optical properties. As the sub-images will be turned into a binary image, the discontinuity of the boundary between different sub-images does not impact the image segmentation task.
3.2 Generating binary cell masks